OpenStudy (anonymous):

If the 35th term of an A.P is 30, prove that the sum of its 69 terms will be 2070

OpenStudy (raden):

actually, just multiplying 30 by 69 :V

hartnn (hartnn):

use the n'th term formula \(\Large a_n = a_1 +(n-1)d\) plug in n=35 \(a_{35} = 30\) you will get an equation in a1 and 'd' Then use the sum formula, \(\Large S_n = (n/2)(2a_1 +(n-1)d)\) and use the equation you got before.

hartnn (hartnn):

yes, it would be same as multiplying 30 by 69. if you follow that process, you'll know why :)

OpenStudy (raden):

dont hoped, i was just guessing :)

hartnn (hartnn):

that happened because n-1 = 34 and 2(n-1) = 68 which is also 69-1

OpenStudy (raden):

ohhh yes, i checked it that like i thought :P S69 = 69/2 (2a1 + 68) = 69/2 * 2 * (a1 + 34) = 69 * 30 = 2070

hartnn (hartnn):

\(\huge \checkmark \)

hartnn (hartnn):

Rakesh, ask if any doubts? :)

OpenStudy (anonymous):

69*30 HOW IT COMES

hartnn (hartnn):

did you follow the steps in my first comment ? where are you stuck ?

OpenStudy (anonymous):

a1+34 = 30 cant get

hartnn (hartnn):

you missed the 'd' \(\Large 30 = a_1 +34 d \quad \to (1)\) thats your 1st equation

hartnn (hartnn):

now use the sum formula

OpenStudy (anonymous):

S69 = 69/2 (2a1 + 68) = 69/2 * 2 * (a1 + 34) = 69 * 30 = 2070 wanna understand this only

hartnn (hartnn):

did you try plugging in values in the sum formula by yourself ?

hartnn (hartnn):

forget that you ever saw that line

OpenStudy (anonymous):

ok

hartnn (hartnn):

could you complete the problem ? \(\Large S_{69} = ... ?\)

OpenStudy (anonymous):

I UNDERSTAND D FORMULA

OpenStudy (anonymous):

only d problem with that 30

OpenStudy (anonymous):

s69=69/2(2a+68), = 69/2*2(a+34), 69*(a+34)

hartnn (hartnn):

you again missed the 'd' s69=69/2(2a+68\(\Large d\)), = 69/2*2(a+34\(\Large d\)), =69*(a+34\(\Large d\)) and using equation (1), = 69*30 =2070